# Data-driven computation of adjoint sensitivities without adjoint solvers: An application to thermoacoustics

### v1

@@ VI.3 Computation of adjoint sensitivities
Figure 4: Adjoint sensitivities of the true system and adjoint sensitivities inferred by the thermoacoustic Echo State Network over an integration time of 100 time units. Top row, sensitivity of the time-averaged acoustic energy to heat release strength $\beta$, $dJ/d\beta$ at fixed (a) $\tau=0.07$, (b) $\tau=0.12$, (c) $\tau=0.22$, (d) $\tau=0.32$. Bottom row, sensitivity of the time-averaged acoustic energy to time-delay $\tau$, $dJ/d\tau$ at fixed (e) $\beta=1.25$, (f), $\beta=2.5$, (g)$\beta=3.75$, (h) $\beta=4.5$. The mean of the sensitivity of an ensemble of 5 realisations of the ESN is shown, one standard deviation from the mean is plotted in lighter colour.
[…]
Fig. 4 also shows that the T-ESN infers the existence of bifurcations to fixed point solutions. The fixed point of the Rijke tube is the zero vector, i.e., the mean flow, which means that the sensitivity is zero for the fixed point regimes, e.g., at fixed $\tau=0.22$ (Fig. 4(c)) for $\beta<0.45$. Albeit the parameter value at which the bifurcation occurs might have a small error compared to the true system. For example, at fixed $\tau=0.22$ the Hopf bifurcation occurs at $\beta=0.45$, whereas 4 out of 5 realisations of the T-ESN predict it as $\beta\approx 0.35$, and the remaining realisation predicts it as $\beta\approx 0.30$.
@@ VII Gradient-based optimization
Figure 6: Gradient-based optimization of time-averaged acoustic energy marked as Path 3 in Fig. 5. Shown (a) iteration 0; $(\beta=4.0,\tau=0.25)$, (b) iteration 2; $(\beta=3.57,\tau=0.17)$, (c) iteration 4; $(\beta=2.85,\tau=0.1)$, and (d) iteration 8; $(\beta=1.64,\tau=0.07)$. From left to right; the first two pressure modes in the phase space, short-term timeseries prediction of the acoustic energy, and long-term amplitude spectrum of the acoustic energy. The optimization follows a path along which the amplitude of the oscillations and the acoustic energy decreases, while the ESN accurately predicts the changing dynamics.
@@
We apply the inferred adjoint sensitivities to a gradient-based optimization framework. One realisation of the T-ESN suffices for the computation of the gradients and optimization for all parameters. In this work, the goal of the optimization is to determine the set of parameters $(\beta,\tau)$ that minimizes the time-averaged acoustic energy (20). We restrict the domain to $\beta=[0.5,5.5],\;\tau=[0.05,0.35]$, which prevents the parameters to take negative (non-physical) values, and the system to bifurcate to quasiperiodic or chaotic regimes. We wish to avoid this during gradient-based optimization because first, the acoustic energy tends to display discontinuous jumps at the parameter values where the system bifurcates [6], which can lead to undefined gradients between different nonlinear regimes. Second, while the adjoint sensitivity can be computed for limit-cycle and quasiperiodic regimes for arbitrary integration times, for chaotic regimes the integration time is limited before the adjoint system becomes unstable, and the sensitivities diverge [50, 6]. Another consideration is the respective orders of magnitudes of the parameters and the sensitivities to them. From Fig. 4, $\beta$ and $\tau$ as well as $dJ/d\beta$ and $dJ/d\beta$ have different orders of magnitudes, i.e., $\beta\sim O(1),\;\tau\sim O(10^{-1})$. The difference in the orders of magnitudes requires the parameters and their respective gradients to be normalized, such that $\breve{\beta}=\beta/\bar{\beta},\;\breve{\tau}=\tau/\bar{\tau},\;dJ/d\breve{\beta}=\bar{\beta}dJ/d\beta,\;dJ/d\breve{\tau}=\bar{\tau}dJ/d\tau$, where $\bar{(\cdot)}$ denotes the mean of the parameter in the domain considered [6]. The optimization is performed with the steepest-descent method using the normalized gradients with a step size of 0.2. (Since $\tau$ can only take discrete values, we round it according to the time-step size of the ESN before the next step of the optimization.) The optimization stops when (i) the acoustic energy is near zero, $J_{i}<10^{-4}$, or (ii) the relative change in acoustic energy is below a threshold, $(J_{i}-J_{i-1})/(J_{i-1})<10^{-4}$ where subscript $i$ denotes the $i^{th}$ iteration step. Iterations of optimization starting from different initial $(\beta,\tau)$ are shown in Fig. 5. The optimal $(\beta,\tau)$ values are found to be in the fixed point regimes, i.e., when the instability is suppressed. For the Rijke tube, the fixed point is the zero vector (mean flow), and as such the acoustic energy, i.e., the acoustic energy of the fluctuations from the mean flow, is equal to zero at these values. This result shows the significance of the ability of the ESN to infer fixed points. The process of optimization is illustrated for Path 3 in Fig 6. The optimization follows a path where the amplitude of the oscillations and so the acoustic energy decreases, for which the ESN accurately predicts the changing dynamics. Similar behaviour is observed for other paths. An exceptional case is Path 4, which passes through a saddle in the landscape of the objective function near $(\beta=4.5,\tau=0.15)$. The change in the sign of the gradient, as shown in Fig. 4 in the sensitivity plot for $\beta=4.5$ around $\tau=0.13$, causes the “zigzagging” path in this region.
VIII Robutsness to noise
@@
If the data is acquired from real-life experiments, the measurements may be corrupted by noise. This is the case when the system equations are unknown and we wish to learn sensitivities from experimental data. To test the performance of the T-ESN under such a scenario, we consider zero-mean additive Gaussian noise with a standard deviation that is equal to 5% of the standard deviation of the true data (this corresponds to 26 dB signal-to-noise ratio). We repeat the training on the noisy data using the hyperparameters chosen before (Tab. 1) and three Tikhonov coefficients, $\lambda=10^{-6}$ as before, $\lambda=10^{-5}$ and $\lambda=10^{-3}$ to observe the effect of increasing regularization. We show the performance of closed-loop prediction on a parameter regime $(\beta=2.0,\tau=0.25)$, which is in the training set, in Fig. 7. The T-ESN successfully learns the limit cycle attractor from noisy data and does not become unstable when the training data is noisy. We compute the adjoint sensitivities for the same test cases in Sec.VI.3, these are shown in Fig. 8. The sensitivity computations are affected by the noisy training especially in small $\tau$ regimes ($\tau\leq 0.1$) that are unseen during training. These errors are alleviated by increasing the regularization. However, using a very large regularization might decrease the performance, for example in higher $\tau$ regimes ($\tau\geq 0.25$) as shown in Fig. 7.
@@
Figure 7: Thermoacoustic Echo State Network learns the true limit cycle attractors of the original system from noisy data. Short-term closed-loop prediction on the training data from $(\beta=2.0,\tau=0.25)$ regime is shown for the first pressure mode as timeseries (a) and for the first and second pressure modes in the phase space (b). The amplitude spectrum (c) is computed for the same time window as the training data (8 time units) for consistency in frequency resolution and shown in logarithmic scale. Short-term prediction, probability distribution function (PDF), and amplitude spectrum of the acoustic energy are shown as in Fig. 3.
Figure 8: Adjoint sensitivities of the true system and adjoint sensitivities inferred by the thermoacoustic Echo State Network when the training is done on noisy data. Top row, sensitivity of the time-averaged acoustic energy to heat release strength $\beta$, $dJ/d\beta$ at fixed (a) $\tau=0.07$, (b) $\tau=0.12$, (c) $\tau=0.22$, (d) $\tau=0.32$. Bottom row, sensitivity of the time-averaged acoustic energy to time-delay $\tau$, $dJ/d\tau$ at fixed (e) $\beta=1.25$, (f), $\beta=2.5$, (g)$\beta=3.75$, (h) $\beta=4.5$. The mean of the sensitivity of an ensemble of 5 realisations of the ESN is shown, one standard deviation from the mean is plotted in lighter colour. Increasing Tikhonov coefficient regularizes the training.
@@ IX Bifurcations to nonlinear regimes
Refer to captionRefer to caption
Figure 9: Thermoacoustic Echo State Network captures bifurcations to period-doubling and quasi-periodic regimes. Prediction, statistics, and amplitude spectrum of (a-f) an unseen limit cycle regime where period-doubling occurs ($\beta=7.5,\;\tau=0.3$), (g-l) an unseen quasi-periodic regime ($\beta=6.1,\;\tau=0.2$)
Refer to captionRefer to caption
@@
Figure 11: Bifurcation diagrams obtained from the true system (top row), thermoacoustic ESN trained on $\beta=[1,5]$ (middle row), and trained on $\beta=[6,8]$ (bottom row). The peaks of the first pressure mode $\mu_{1}$ and the acoustic energy are plotted. Left half (a-f) shows varying $\beta$ for fixed $\tau=0.22$, and right half (g-l) shows varying $\tau$ for fixed beta = 7.6.
@@
In Sec. VI, we discuss that the ESN predicts bifurcations from limit cycle oscillations to fixed points when the heat release strength $\beta$ is decreased. On the other hand, if $\beta$ is increased, the thermoacoustic system displays bifurcations from limit cycle oscillations to chaos via period-doubling and quasiperiodic routes [6]. In this section, we investigate whether the ESN can also predict these bifurcations and attractors. We train and validate a separate T-ESN on data generated by larger values $\beta=\{6.0,6.5,7.0,7.5,8.0\}$ (for brevity, we refer to this dataset as $\beta_{train}=[6.0-8.0]$). The reservoir size and connectivity are 1200 and 20, respectively as before. The optimal hyperparameters found for this case are provided in Table 2 along with the hyperparameters for the previous case when the training was done with data from $\beta=\{1.0,2.0,3.0,4.0,5.0\}$ (for brevity, we refer to this dataset as $\beta_{train}=[1.0-5.0]$). The faster changing dynamics, e.g., chaotic behaviour, in the larger $\beta$ regimes might be responsible for the increase in the spectral radius and leak rate.
@@
In Fig. 9(a-f) we show a limit cycle regime where period-doubling occurs ($\beta=7.5,\;\tau=0.3$), and in Fig. 9(g-l) a quasi-periodic regime ($\beta=6.1,\;\tau=0.2$). Chaotic regimes are characterized by the time at which two nearby trajectories diverge exponentially. This characteristic timescale is known as the Lyapunov time (LT) and is an indicator of the predictability of a system. We examine two chaotic thermoacoustic regimes with different Lyapunov times; in Fig. 10(a-f) a regime ($\beta=7.6,\;\tau=0.22$) with a Lyapunov time of 8.5 time units and in Fig. 10(g-l) a regime ($\beta=8.7,\;\tau=0.23$) with a Lyapunov time of 3.9 time units. The ability of the T-ESNs trained on different $\beta$ regimes to predict bifurcations is further shown in Fig. 11 in the bifurcation diagrams for fixed $\tau=0.22$ and $\beta=7.6$ cases. The ESN trained on $\beta_{train}=[1.0-5.0]$ can qualitatively predict the bifurcations and capture the statistics and dominant frequencies correctly, especially for the non-chaotic regimes. This ability of the parameter-aware ESN to predict bifurcations to quasi-periodic or chaotic regimes at unseen parameter values is in itself not an obvious result as the network only had access to limit cycle regimes at lower $\beta$ values for training and validation. For example, when we increase $\beta$ and a bifurcation occurs, the acoustic energy increases with a discontinuous jump, which indicates that the amplitude of the oscillations abruptly change. Therefore, the prediction of these regimes becomes an extrapolation problem because the range of data we are interested in is different from the range of data the network was trained on. In this case, incorporating physical knowledge via the T-ESN architecture enables extrapolation performance. Training the T-ESN on $\beta_{train}=[6.0-8.0]$, which encompasses the regimes that we are testing, increases accuracy of both short- and long-term predictions. The bifurcation diagrams confirm that in the higher $\beta$ regimes, the T-ESN trained on $\beta_{train}=[1.0-5.0]$ can qualitatively capture the bifurcations to quasiperiodic and chaotic regimes, however is not as quantitatively accurate as the T-ESN trained on $\beta_{train}=[6.0-8.0]$. (The reverse is true for the lower $\beta$ regimes.) When $\beta$ is further increased beyond $\beta=10.0$, the quantitatively accurate performance of the T-ESN trained on $\beta_{train}=[6.0-8.0]$ also starts diminishing.
@@
In chaotic systems, due to the positive Lyapunov exponent, the adjoint system is unstable and the adjoint sensitivities diverge [50, 6]. This means that we can only integrate adjoint equations for short time windows. Using the T-ESN trained on $\beta_{train}=[6.0-8.0]$, we compute the adjoint sensitivities of short-time trajectories for the ($\beta=7.6,\;\tau=0.22$) regime. Figures 12(a,b) show the evolution of the adjoint sensitivities with integration time for one trajectory for an ensemble of 5 T-ESNs. For short integration times, the T-ESN predicts the chaotic sensitivities close to the original system, but diverges as the integration time increases due to the chaotic nature of the system. One approach to estimate the sensitivity of chaotic systems is by taking the mean of an ensemble trajectories [50]. In Figs. 12(c,d) and (e,f), we show the probability distributions and the mean of the adjoint sensitivities for 2000 trajectories for integration times of 0.5 LT and 1 LT, respectively. In order to obtain the PDFs, we let the T-ESN run autonomously without any washout data from the original system. For an integration time of 0.5 LT, we can recover the PDF as well as the mean, while for 1.0 LT, low statistical moments are matched.

### v2

@@ VI.3 Computation of adjoint sensitivities
Figure 4: Thermoacoustic Echo State Network infers sensitivity of the time-averaged acoustic energy, i.e., objective $\mathcal{J}$, to heat release strength $\beta$. Shown $d\mathcal{J}/d\beta$ at fixed (a) $\tau=0.07$, (b) $\tau=0.12$, (c) $\tau=0.22$, (d) $\tau=0.32$. The mean of the sensitivity of an ensemble of 5 realisations of the ESN is shown, one standard deviation from the mean is plotted in lighter colour.
Figure 5: Thermoacoustic Echo State Network infers sensitivity of the time-averaged acoustic energy, i.e., objective $\mathcal{J}$, to time-delay $\tau$. Shown $d\mathcal{J}/d\tau$ at fixed (a) $\beta=1.25$, (b) $\beta=2.5$, (c) $\beta=3.75$, (d) $\beta=4.5$. The mean of the sensitivity of an ensemble of 5 realisations of the ESN is shown, one standard deviation from the mean is plotted in lighter colour.
We compare the adjoint sensitivities computed by the T-ESN following the procedure described in Sec. III with the adjoint sensitivities obtained from the original system described by Eq. (22). We refer to the latter as the ground truth, which is computed by deriving the Jacobian of the ODEs (22) and integrating the adjoint equations (5). The sensitivities are shown for varying $\beta$ and $\tau$ values in Figs. 4 and 5, respectively. For ease of results, we show the sensitivity to $\beta$ when $\beta$ is varied and $\tau$ is fixed, and vice versa. In the washout stage, we feed the initial condition of zeros except $\eta_{1}(t_{0})=1.5$, assuming that no data is available from the original system for the washout. This value of $\eta_{1}(t_{0})$ makes sure that the trajectories converge to the stable limit cycle and not the stable fixed point in the multistable regimes. After the washout and the transient, we solve the adjoint equations around a trajectory of 100 time units, for which we find that the adjoint sensitivity converges to a steady-state value. The learned adjoint sensitivities match closely with the ground truth, while the results are also robust to different realisations of the network. Even though the training data was generated from discrete values of $(\beta,\tau)$ on a grid (Fig. 2), we are able to obtain sensitivity information from a range of $(\beta,\tau)$ values that are unseen during the training using a single instance of the T-ESN. The smoothness of the sensitivity curve also indicates that the dynamics of the T-ESN varies smoothly with the parameters. The accuracy decreases as we move away from the training regimes, e.g., beyond $\beta=5.0$ in the $d\mathcal{J}/d\beta$ plot at $\tau=0.32$ (Fig. 4 (d)), but the ESN captures the trend of the sensitivities.
Figs. 4 and 5 also show that the T-ESN infers the existence of bifurcations to fixed point solutions. The fixed point of the Rijke tube is the zero vector, i.e., the mean flow, which means that the sensitivity is zero for the fixed point regimes, e.g., at fixed $\tau=0.22$ (Fig. 4 (c)) for $\beta<0.45$. Albeit the parameter value at which the bifurcation occurs might have a small error compared to the true system. For example, at fixed $\tau=0.22$ the Hopf bifurcation occurs at $\beta=0.45$, whereas 4 out of 5 realisations of the T-ESN predict it as $\beta\approx 0.35$, and the remaining realisation predicts it as $\beta\approx 0.30$.
@@ VII Gradient-based optimization
Figure 7: Gradient-based optimization of time-averaged acoustic energy marked as Path 3 in Fig. 6. Shown (a-c) iteration 0; $(\beta=4.0,\tau=0.25)$, (d-f) iteration 2; $(\beta=3.57,\tau=0.17)$, (g-i) iteration 4; $(\beta=2.85,\tau=0.1)$, and (j-l) iteration 8; $(\beta=1.64,\tau=0.07)$. From left to right; the first two pressure modes in the phase space, short-term timeseries prediction of the acoustic energy, and long-term amplitude spectrum of the acoustic energy. The optimization follows a path along which the amplitude of the oscillations and the acoustic energy decreases, while the thermoacoustic Echo State Network accurately predicts the changing dynamics.
@@
We apply the inferred adjoint sensitivities to a gradient-based optimization framework. One realisation of the T-ESN suffices for the computation of the gradients and optimization for all parameters. In this work, the goal of the optimization is to determine the set of parameters $(\beta,\tau)$ that minimizes the time-averaged acoustic energy (25). We restrict the domain to $\beta=[0.5,5.5],\;\tau=[0.05,0.35]$, which prevents the parameters to take negative (non-physical) values, and the system to bifurcate to quasiperiodic or chaotic regimes. We wish to avoid this during gradient-based optimization because first, the acoustic energy tends to display discontinuous jumps at the parameter values where the system bifurcates [6], which can lead to undefined gradients between different nonlinear regimes. Second, while the adjoint sensitivity can be computed for limit-cycle and quasiperiodic regimes for arbitrary integration times, for chaotic regimes the integration time is limited before the adjoint system becomes unstable, and the sensitivities diverge [48, 6]. We deal with the sensitivity of the chaotic regimes separately in Sec. IX. Another consideration is the respective orders of magnitudes of the parameters and the sensitivities to them. From Figs. 4 and 5, $\beta$ and $\tau$ as well as $d\mathcal{J}/d\beta$ and $d\mathcal{J}/d\beta$ have different orders of magnitudes, i.e., $\beta\sim O(1),\;\tau\sim O(10^{-1})$. The difference in the orders of magnitudes requires the parameters and their respective gradients to be normalized, such that $\breve{\beta}=\beta/\bar{\beta},\;\breve{\tau}=\tau/\bar{\tau},\;d\mathcal{J}/d\breve{\beta}=\bar{\beta}d\mathcal{J}/d\beta,\;d\mathcal{J}/d\breve{\tau}=\bar{\tau}d\mathcal{J}/d\tau$, where $\bar{(\cdot)}$ denotes the mean of the parameter in the domain considered [6]. The optimization is performed with the steepest-descent method using the normalized gradients with a step size of 0.2. (Since $\tau$ can only take discrete values, we round it according to the time-step size of the ESN before the next step of the optimization.) The optimization stops when (i) the acoustic energy is near zero, $\mathcal{J}_{i}<10^{-4}$, or (ii) the relative change in acoustic energy is below a threshold, $(\mathcal{J}_{i}-\mathcal{J}_{i-1})/(\mathcal{J}_{i-1})<10^{-4}$ where subscript $i$ denotes the $i^{th}$ iteration step. Iterations of optimization starting from different initial $(\beta,\tau)$ are shown in Fig. 6. The optimal $(\beta,\tau)$ values are found to be in the fixed point regimes, i.e., when the instability is suppressed. For the Rijke tube, the fixed point is the zero vector (mean flow), and as such the acoustic energy, i.e., the acoustic energy of the fluctuations from the mean flow, is equal to zero at these values. This result shows the significance of the ability of the ESN to infer fixed points. The process of optimization is illustrated for Path 3 in Fig 7. The optimization follows a path where the amplitude of the oscillations and so the acoustic energy decreases, for which the ESN accurately predicts the changing dynamics. Similar behaviour is observed for other paths. An exceptional case is Path 4, which passes through a saddle in the landscape of the objective function near $(\beta=4.5,\tau=0.15)$. The change in the sign of the gradient, as shown in Fig. 5 (d) in the sensitivity plot for $\beta=4.5$ around $\tau=0.13$, causes the “zigzagging” path in this region.
VIII Robustness to noise
If the data is acquired from real-life experiments, the measurements may be corrupted by noise. This is the case when the system equations are unknown and we wish to learn sensitivities from experimental data. To test the performance of the T-ESN under such a scenario, we consider zero-mean additive Gaussian noise with a standard deviation that is equal to 2%, 5%, and 10% of the standard deviation of the true data (these noise levels correspond to 33, 26, and 20 dB signal-to-noise ratios, respectively). We repeat the training on the noisy data using the hyperparameters chosen before (Tab. 1) and investigate the effect of regularization. The regularization parameter, $\lambda$ can be determined by evaluating the validation error during hyperparameter selection. Fig. 8 shows the statistics (mean and one standard deviation) of the validation error of the T-ESN for different levels of noise. The validation error is given by the mean relative $\ell_{2}$ error (32) as described in Sec. VI.1. (In the noisy case, we increase the number of evaluated trajectories to 20 as this helps with the covergence of the error statistics.) To obtain the error statistics, we repeat this calculation for 20 different realisations of noise applied to the training and validation data. Because the validation data contains noise, the irreducible error is the noise level. We find that there is a range of $\lambda$ values that results in a similarly low validation error, and this optimal range of regularization increases with increasing levels of noise.
Figure 8: Selection of the regularization parameter when the data is noisy. Shown is the validation error of thermoacoustic Echo State Network on data with different levels of noise (indicated in the legend) when the training is performed with increasing regularization. The error statistics for 20 different realisations of noise are plotted.
@@ VIII Robustness to noise
For the 5% noise level case, we show the performance of closed-loop prediction on a parameter regime $(\beta=2.0,\tau=0.25)$, which is in the training set, in Fig. 9. The T-ESN successfully learns the limit cycle attractor from noisy data and does not become unstable when the training data is noisy. We compute the adjoint sensitivities for the same test cases in Sec.VI.3, these are shown in Figs. 10 and 11 for the heat release strength $\beta$ and time delay $\tau$, respectively. The sensitivity computations are affected by the noisy training especially in small $\tau$ regimes ($\tau\leq 0.1$) that are unseen during training, e.g., Fig. 10 (a,b). These errors are alleviated by increasing the regularization. However, using a very large regularization might decrease the performance, for example in higher $\tau$ regimes ($\tau\geq 0.25$) as shown in Figs. 9 and 11(b).
@@
Figure 9: Thermoacoustic Echo State Network learns the true limit cycle attractors of the original system from noisy data. Short-term closed-loop prediction on the training data from $(\beta=2.0,\tau=0.25)$ regime is shown for the first pressure mode as timeseries (a) and for the first and second pressure modes in the phase space (b). The amplitude spectrum (c) is computed for the same time window as the training data (8 time units) for consistency in frequency resolution and shown in logarithmic scale. Short-term prediction (d), probability distribution function (PDF) (e), and amplitude spectrum (f) of the acoustic energy are shown as in Fig. 3.
Figure 10: Regularization with an optimal Tikhonov coefficient $\lambda=10^{-3}$ improves inference of sensitivity to heat release strength in case of noise level 5% (See Fig. 8). Shown $d\mathcal{J}/d\beta$ at fixed (a) $\tau=0.07$, (b) $\tau=0.12$, (c) $\tau=0.22$, (d) $\tau=0.32$. The mean of the sensitivity of an ensemble of 5 realisations of the ESN is shown for different regularization coefficients, one standard deviation from the mean is plotted in lighter colour.
Figure 11: Regularization with an optimal Tikhonov coefficient $\lambda=10^{-3}$ improves inference of sensitivity to time delay $\tau$ in case of noise level 5% (See Fig. 8). Shown $d\mathcal{J}/d\tau$ at fixed (a) $\beta=1.25$, (b) $\beta=2.5$, (c) $\beta=3.75$, (d) $\beta=4.5$. The mean of the sensitivity of an ensemble of 5 realisations of the ESN is shown for different regularization coefficients, one standard deviation from the mean is plotted in lighter colour.
@@ IX Bifurcations to nonlinear regimes
Figure 12: Thermoacoustic Echo State Network captures bifurcations to period-doubling and quasi-periodic regimes. Short-term prediction, statistics, and amplitude spectrum of (a-f) an unseen limit cycle regime where period-doubling occurs ($\beta=7.5,\;\tau=0.3$), (g-l) an unseen quasi-periodic regime ($\beta=6.1,\;\tau=0.2$)
@@
Figure 14: Bifurcation diagrams obtained from the true system (top row), thermoacoustic Echo State Network trained on $\beta_{train}=[1,5]$ (middle row), and trained on $\beta_{train}=[6,8]$ (bottom row). The peaks of the acoustic energy are plotted. Left half (a,b,c) shows varying $\beta$ for fixed $\tau=0.22$, and right half (d,e,f) shows varying $\tau$ for fixed $\beta=7.6$.
@@
In Sec. VI, we discuss that the ESN predicts bifurcations from limit cycle oscillations to fixed points when the heat release strength $\beta$ is decreased. On the other hand, if $\beta$ is increased, the thermoacoustic system displays bifurcations from limit cycle oscillations to chaos via period-doubling and quasiperiodic routes [6]. In this section, we investigate whether the ESN can also predict these bifurcations and attractors. We train and validate a separate T-ESN on data generated by larger values $\beta=\{6.0,6.5,7.0,7.5,8.0\}$ (for brevity, we refer to this dataset as $\beta_{train}=[6.0-8.0]$). The reservoir size and connectivity are 1200 and 20, respectively as before. The optimal hyperparameters found for this case are provided in Table 2 along with the hyperparameters for the previous case when the training was performed with data from $\beta=\{1.0,2.0,3.0,4.0,5.0\}$ (for brevity, we refer to this dataset as $\beta_{train}=[1.0-5.0]$). The faster changing dynamics, e.g., chaotic behaviour, in the larger $\beta$ regimes might be responsible for the increase in the spectral radius and leak rate.
@@
We examine the T-ESN’s prediction performance across different nonlinear regimes. First, in Fig. 12 (a-c) we show a limit cycle regime in which period-doubling occurs ($\beta=7.5,\;\tau=0.3$), and in Fig. 12 (d-f) a quasi-periodic regime ($\beta=6.1,\;\tau=0.2$). Second, we show two chaotic thermoacoustic regimes with different Lyapunov times; in Fig. 13 (a-f) a regime ($\beta=7.6,\;\tau=0.22$) with a Lyapunov time of 8.5 time units and in Fig. 13 (g-l) a regime ($\beta=8.7,\;\tau=0.23$) with a Lyapunov time of 3.9 time units. Lyapunov time (LT) is the characteristic timescale, by which two nearby trajectories diverge exponentially in a chaotic system and is an indicator of the predictability of a system. The ESN trained on $\beta_{train}=[1.0-5.0]$ can qualitatively predict the bifurcations and capture the statistics and dominant frequencies correctly, especially for the non-chaotic regimes. Training the T-ESN on $\beta_{train}=[6.0-8.0]$, which encompasses the regimes that we are testing, increases accuracy of both short- and long-term predictions.
The ability of the T-ESNs trained on different $\beta$ regimes to predict bifurcations is further shown in Fig. 14 in the bifurcation diagrams for fixed $\tau=0.22$ (a-c) and $\beta=7.6$ (d-f) cases. This ability of the parameter-aware ESN to predict bifurcations to quasiperiodic or chaotic regimes at unseen parameter values is in itself not an obvious result as the network only had access to limit cycle regimes at lower $\beta$ values for training and validation. When we increase $\beta$ and a bifurcation occurs, the acoustic energy increases with a discontinuous jump, which indicates that the amplitude of the oscillations abruptly change. Therefore, the prediction of these regimes becomes an extrapolation problem because the range of data we are interested in is different from the range of data the network was trained on. In this case, incorporating physical knowledge via the T-ESN architecture enables extrapolation performance. The bifurcation diagrams confirm that in the higher $\beta$ regimes, the T-ESN trained on $\beta_{train}=[1.0-5.0]$ can qualitatively capture the bifurcations to quasiperiodic and chaotic regimes, however is not as quantitatively accurate as the T-ESN trained on $\beta_{train}=[6.0-8.0]$. (The reverse is true for lower $\beta$ regimes.) When $\beta$ is further increased beyond $\beta=10.0$, the quantitatively accurate performance of the T-ESN trained on $\beta_{train}=[6.0-8.0]$ also starts diminishing.
@@
In chaotic systems, due to the positive Lyapunov exponent, the adjoint system is unstable and the adjoint sensitivities diverge [48, 6]. The ESN learns and replicates the underlying chaotic dynamics of the system under investigation [e.g., 18]. Because the ESN is a dynamical system (10), the adjoint sensitivity in chaotic regimes is unstable in the long-term [6]. This means that we can only integrate adjoint equations for short time windows. Using the T-ESN trained on $\beta_{train}=[6.0-8.0]$, we compute the adjoint sensitivities of short-time trajectories for the ($\beta=7.6,\;\tau=0.22$) regime (shown in Fig. 13 (a-c)). Figs. 15 (a,b) show the evolution of the adjoint sensitivities with integration time for one trajectory for an ensemble of 5 T-ESNs. For short integration times, the T-ESN predicts the chaotic sensitivities close to the original system, but diverges as the integration time increases due to the chaotic nature of the system. One approach to estimate the sensitivity of chaotic systems is by taking the mean of an ensemble trajectories [48]. In Figs. 15 (c,d) and (e,f), we show the probability distributions and the mean of the adjoint sensitivities for 2000 trajectories for integration times of 0.5 LT and 1 LT, respectively. In order to obtain the PDFs, we let the T-ESN run autonomously without any washout data from the original system. For an integration time of 0.5 LT, we can recover the PDF as well as the mean, while for 1.0 LT, low statistical moments are matched.